OpenStudy (anonymous):

QUIZ:

OpenStudy (anonymous):

Solve: \[\frac{ dy }{ dx } = \frac{ 2\cos 2x }{ 3+2y }\] with y(0) =-1 (i) For what values of 𝑥 > 0 does the solution exist? (ii) For what value of 𝑥 > 0 is 𝑦(𝑥) maximum?

OpenStudy (anonymous):

@PaxPolaris any idea??

OpenStudy (paxpolaris):

fill in the blanks in you questions ...

OpenStudy (paxpolaris):

\[(3+2y)dy=(2\cos(2x))dx\]

OpenStudy (paxpolaris):

(i) For what values of 𝑥___ > 0 does the solution exist? (ii) For what value of 𝑥___ > 0 is 𝑦(___) maximum? can i assume it's x, x and y?

OpenStudy (anonymous):

yes of course., please

OpenStudy (paxpolaris):

integrate both sides:\[\implies \int\limits (3+2y)dy=\int\limits 2\cos(2x)dx\]

OpenStudy (anonymous):

𝑦2 + 3𝑦 + 2 = sin 2𝑥 right??? then??

OpenStudy (paxpolaris):

\[\large y^2+3y+2=\sin(2x)\] ?

OpenStudy (anonymous):

because y(0) = -1, so \[\int\limits_{-1}^{y} 3 + 2y dy = \int\limits_{0}^{x} 2 \cos 2x dx\]

OpenStudy (anonymous):

solve it 3+2y(-2xin2x.co2x) - (2cos2x)2 divided by (3+2y)^2 u ll get the ans

OpenStudy (anonymous):

the above equation came by differentiating the equation

OpenStudy (paxpolaris):

\[\Large \implies y= {-3 \pm \sqrt{9-4(2-\sin(2x)}\over 2}={-3 \pm \sqrt{1+4\sin(2x)}\over 2}\]

OpenStudy (paxpolaris):

(i) y has solution when:\[\Large 1+4\sin(2x) \ge 0\]

OpenStudy (paxpolaris):

ps. i just used the quadratic formula

OpenStudy (anonymous):

then??

OpenStudy (paxpolaris):

 you could just use wolfram alpha for the answer